OpenStudy (anonymous):

Solve sin(4x)=-2sin(x) ...sin(2x+2x)+2sin(x)=0

OpenStudy (bahrom7893):

Look up some double angle formulas.. What's Sin(2a)?

OpenStudy (bahrom7893):

Sin(4x) = Sin(2*2x) = 2Sin(2x)Cos(2x) = -2Sin(x)

OpenStudy (bahrom7893):

\[2(2Sin(x)*Cos(x))*(Cos^2(x) - Sin^2(x))=-2Sin(x)\]

OpenStudy (bahrom7893):

Dividing both sides by \[2Sin(x)\]we get:\[2Cos(x)(Cos^2(x)-Sin^2(x))=-1\]

OpenStudy (bahrom7893):

Aandd now I'm stuck.. im sure this is just some algebra and more simplification... hang on..

OpenStudy (bahrom7893):

Well actually -1 = -(Sin^2 + Cos^2)

OpenStudy (bahrom7893):

uhm yea that doesn't really help us anyway..

OpenStudy (anonymous):

Looking for the solutions, which you should be able to get from the factors. Th answers should be 0, (pi/2), (pi), (3pi/2)

OpenStudy (bahrom7893):

Wolfram is doing some crazy substitutions... i dont think they're necessary..

OpenStudy (bahrom7893):

@hartnn @Zarkon

hartnn (hartnn):

sin 4x = -2sin x 2 sin 2x cos 2x = -2 sin x 4 sin x cos x (2 cos^2x -1) + 2 sinx =0 factoring out 2 sin x 2 sin x [2 cos x (2cos^2x -1)+ 2 ] = 0 which gives us sin x =0 OR 2 cos^3 x - cos x + 1 = 0 if we let cos x = p we have a cubic function in p 2p^3 -p +1 = 0 solve this to get p = cos x = .... , ... , ...